OpenStudy (anonymous):

Find all values of x in the interval (0,2Pi) that satisfy equation 2sinxcosx=cosx

OpenStudy (experimentx):

2sinxcosx = sin(2x) = cos(pi/2 -x) now cos(pi/2 -x) = cos(x) x = pi/4

OpenStudy (anonymous):

thank you